A firm operating in a competitive industry has it cost function as TC =100+4Q-10Q²+⅓Q² if the market price P=30


Require

Find the profit maximizing unit of outline and the total profit at the level of output ratio


1
Expert's answer
2021-11-07T19:46:11-0500

The firm will maximize profit by producing at a point where the marginal cost is equal to marginal revenue.

"TC =100+4Q-10Q^²+\frac{1}{3}Q^3"

"MC=4-20Q+Q^2"

"TR=P\times Q\\=30Q"

"MR=30"


"MR=MC\\30=4-20Q+Q^2\\Q^2-20Q-26=0\\Q=21.225\\\approx21"


"profit=TR-TC\\=[30\times 21]-[100+4(21)-10(21)^2+\frac{1}{3}(21)^3]=1,769"


so the maximizing units of output is 21 and the profit is 1,769